I need help with this questions, it's all in 1
I'm confused
Two Bags contain marbles
bag 1: 16 red, 13 blue
bag 2: 11 red, 9 blue
Rules of the game
Player 1 takes 2 marbles from bag 1, no replacement
Player 2 takes 2 marbles from bag 2, with replacement
a) If player selects a marble from bag 1 what are the odds in favour of selecting a blue marble?
b) What is the probability that player 2 selects 2 blue marbles?
c) What is the probability that player 1 selects 2 blue marbles?
d) If a player selects 2 blue marbles then he/she wins. Is this a fair game, explain
e) If player 1's second marble is blue, what is the probability that his/her first marble was also blue?

Answer by Boreal(15235)

a. 13/29 for one marble. There are 29 marbles and 13 are blue.
b. (9/20)^2=0.2025. With replacement they are independent and one multiplies the initial probability by itself.
c. (13/29)(12/28)=156/812=0.1921. Notice the change in the probability.
d. It is not fair in that the probability is higher for Player 2 to select two blue marbles.
e. With replacement is independent, so the probability of the first marble's being blue is 9/20.